# Interactions across multiple games: cooperation, corruption, and organizational design

### v2

@@ Abstract
It is socially beneficial for teams to cooperate in some situations (“good games”) and not in others (“bad games;” e.g., those that allow for corruption). A team’s cooperation in any given game depends on expectations of cooperation in future iterations of both good and bad games. We identify when sustaining cooperation on good games necessitates cooperation on bad games. We then characterize how a designer should optimally assign workers to teams and teams to tasks that involve varying arrival rates of good and bad games. Our results show how organizational design can be used to promote cooperation while minimizing corruption.
@@ 1 Introduction
Organizations require teamwork: informal cooperation among members. However, team members can also cooperate in ways that are harmful to the organization or society. Importantly, “good” and “bad” forms of cooperation are intertwined. For example, building a cooperative culture among police not only induces productive cooperation, such as aiding other officers in dangerous circumstances, but can also lead them to cooperate in corrupt or destructive ways, such as the “blue wall of silence” (police officers under-reporting misconduct). [^1] See Chevigny (1995). Similarly, some scholars (e.g., Gibson and Singh (2003)) document a “white wall of silence”: doctors not reporting colleagues’ malpractice. If a police officer does not collude with a partner who has taken a bribe or used excessive force, that partner may not back the officer up in a future dangerous situation. [^2] Famously, in 1970, Frank Serpico, then a police detective in the NYPD, exposed police corruption by contributing to a New York Times article. A year later, after being shot and wounded during a drug raid, Serpico alleged that his police partners led him into a dangerous situation to be murdered (Phalen, 2012). Ostracism of whistleblowers is anticipated and widespread: “Officers who report misconduct are ostracized and harassed; become targets of complaints and even physical threats; and are made to fear that they will be left alone on the streets in a time of crisis. This draconian enforcement of the code of silence fuels corruption because it makes corrupt cops feel protected and invulnerable” (Mollen Commission (1994)). Similar interactions between good and bad forms of cooperation arise in a wide variety of organizations, including in government, military, hospitals, consulting, construction, manufacturing, and retail.
@@
Our analysis takes a design perspective. By adjusting team composition and assignments to various tasks, an organizational designer can control the incentives and behavior of the agents involved without observing actions in specific games or even whether a game occurs. This opens a new set of mechanism design problems, in which the dynamic set of tasks faced is controlled by the designer. This design perspective involves several dimensions.
@@
We also allow the assignment of teams to tasks to be dynamic and reactive—i.e., the current task assignment can be a function of past assignments. For example, a team may be rotating between staying on a task with only good games (for some time) and staying on a task with only bad games (for some time). Or a team may be randomly moved between tasks. We also consider the case where the designer observes whether a game occurred (but not the equilibrium play) and can condition future task assignments on that. The space of such “reactive assignments” is large, and so we focus on a special case where cooperation on bad games is easy to sustain. In an optimal assignment, when the designer observes whether a good game occurred, teams are assigned to the task with more good games until a good game occurs, at which point they are reassigned to tasks with mostly bad games for a number of periods before returning. We derive a similar result for the case where the designer does not observe good games: the difference is that teams move on from the task with good games whether or not a good game occurred. As part of this section, we also discuss the continuous limit of our model as time periods and probabilities of a game occurring become vanishingly small—i.e., games arrive according to a Poisson arrival process—and how the previous results are affected.
In Appendix A we provide comparative static results with regard to the payoff structure. If an organization can directly affect payoffs, which is not our focus, then these comparative statics shed light on other strategies for enabling beneficial cooperation. A designer can encourage good cooperation while deterring bad cooperation by altering the benefits of cooperation in good games (for instance, giving bonuses for overall profits) or increasing the temptations of defecting in bad games (for instance, offering whistleblower rewards). Interestingly, these behave quite differently than in standard repeated games, where only the ratio of benefits to temptations matters. With multiple games, benefits from cooperation in a given game have spillovers that enable cooperation in all games, while temptations to deviate affect only one game’s incentives at a time. Thus, the latter is more selective than the former. The effective costs of altering the benefits from cooperation or temptations of defecting also vary. Because bonuses are paid, a policy of giving bonuses for cooperative behavior imposes an actual cost on an organization. In contrast, whistleblower rewards deter collusion and are not dispersed on the equilibrium path.
@@ 2 The Model
Properties 13 are standard: Property 1 states that a player’s payoff is only a function of their own action and the aggregate actions, not their own identity and the identities of others, Property 2 implies that players benefit from cooperation, while Property 3 means each player individually benefits from not cooperating in a single game. Together with Property 4, Property 3 also implies a lower bound on the payoffs a player receives in any continuation, as well as an equilibrium that achieves them. Lastly, Properties 5 and 6 simplify the analysis by ensuring that, on the equilibrium path, asymmetric play is never player-optimal, i.e., in any particular game, there is either full cooperation or full non-cooperation. These properties allow us to characterize the games using just a few parameters described below. Thus, we can focus our attention on interactions across games, rather than the details of dynamics within a game, which can be quite complex absent such structure. [^4] For example, see Olszewski and Safronov (2018). Beyond the dynamic differences, their game also involves private information.
@@
Other games, such as a favor exchange game in which players randomly have chances to either give or receive favors—both good and bad (e.g., lending money vs not reporting a friend’s crime)—also satisfy these properties. Different players move at different points in time, but a similar analysis applies.
@@ 2.2 Player-Optimal Equilibria
As there can be multiple equilibria (the usual folk-theorem arguments apply), we focus on those that are optimal for the players. Given the game’s symmetry and its required properties, we characterize the equilibria that maximize the sum of players’ expected discounted payoffs. [^5] Characterizing such player-optimal equilibria makes sense for two reasons. First, they are extremal—supporting maximal cooperation—and thus define the frontier and so the full set of equilibrium payoffs, regardless of what one believes about equilibrium selection. Second, it may not be robust to presume that a designer can guide players to one equilibrium if all players prefer another.
@@
Visually depicting the different cases of equilibrium cooperation characterized in Lemma 1 is useful. Figure 1 shows which case occurs as a function of the probabilities of good, resp. bad, games occurring times $\delta/(1-\delta)$. Taking Inequality (1) as an equality delineates the region where total cooperation prevails. Cooperation only in some game $g$ precludes total cooperation, so the former region must be southwest of the latter. For cooperation to be sustainable, Inequality (2) must hold; its equality delineates the region with (some) cooperation. Note: there cannot be cooperation only in game $g$ if the temptation to deviate in that game exceeds the deviation temptation in the other game; thus, only one such region exists. The two cases are depicted in the left and right panels of Figure 1. The point labeled $x$ indicates the outcome of the game when teams face a particular pair of probabilities $(p_{G},p_{B})$ and discount factor $\delta$. As shown in the figure, the teams will exhibit total cooperation for these parameters.
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{B}/c_{B}$x\ xregion 2region 1region 3
@@
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{G}/c_{G}$x\ xregion 2region 1region 4
@@
Although the games are independent, their strategic interactions necessitate a holistic analysis to understand players’ actions in one of them. For example, myopically improving the incentives for desirable actions in one type of situation may backfire, generating detrimental behaviors elsewhere. Similarly, collusive behavior may be minimized by eradicating situations where it occurs (decreasing $p_{B}$), reducing its benefits (decreasing $c_{B}$), or incentivizing team members to stop colluding or to report it (increasing $d_{B}$); although all these may serve their intended purpose of reducing collusion, they also can impair productive cooperation in other situations if incentives elsewhere are not adjusted to compensate. The dual of this warning highlights missed opportunities: A myopic design approach may overlook opportunities to reduce bad cooperation by overlooking opportunities to reduce the need for good cooperation. We discuss the implications of changing payoffs in greater detail in Appendix A.
@@ 3 How Cooperation Depends on Team Durability
We model the reshuffling of teams (reducing team durability) as a probability $r$ with which the stochastic game of a team ends after each period. Given the continuum of workers, we assume that reshuffled players never face a past teammate and that there is no spreading of information and that players only condition on histories with their current team. [^7] This greatly simplifies the analysis; see, e.g., Ali and Miller (2016b) for a treatment where such contagion is possible. Our simplification also requires sufficiently many workers so that completely new teams can be formed and anonymity of workers outside their teams (deviations can be punished only by current team members). These conditions are more likely to hold—at least approximately—in large organizations, potentially leading to differences by organization size that we leave for future research. This can also be handled by hiring people part-time or by outsourcing and subcontracting tasks on short-term rotations. Consequently, a reshuffling rate $r$ is equivalent to changing the discount factor $\delta$ to $(1-r)\delta$; essentially, the designer can induce any effective discount factor of at most $\delta$ by varying team durability. Thus, formally, we characterize optimal reshuffling by examining the player optimal equilibria for a given team (with histories only applying to that team) as a function of discount factor $(1-r)\delta$.
@@
The temptation to deviate in good games must be no higher than that in bad games; otherwise, whenever good cooperation is possible, then so is bad. Thus, condition 1 is necessary. Cooperation being sustainable in the good game by itself (condition 2) is also necessary because, for optimal cooperation, there are no strategic interdependencies across the games. Lastly, good games must have a sufficiently high probability, relative to bad games, (condition 3) to make it feasible (possibly by reshuffling) to induce players to cooperate in the good game but not in the bad one: it ensures that there exists an effective discount rate $(1-r)\delta$ such that Equation (1) fails, while (2) holds for the good game. The proposition then follows from a careful application of Lemma 1.
@@ 4 Assigning Teams to Tasks
Next, we consider a second dimension of organizational design affecting cooperation: different teams can be assigned different tasks. The assignment of teams to tasks matters because situations in which teams need to cooperate can occur at different rates across tasks. For example, in a police force, the task of responding to emergencies might create infrequent but important situations where backup is required (good games), whereas the task of traffic policing might present frequent opportunities for accepting small bribes (bad games).
@@
We model this as follows. There is a finite set of tasks $\mathcal{T}$. Each task has its own probability of good and bad games. An assignment for a team, denoted $\nu$, is a probability measure on $\mathcal{T}$ indicating the probability of that team being assigned to different tasks. For example, if there are two tasks, responding to emergencies vs. traffic policing, then $\nu$ indicates the probability that the team is assigned to respond to emergencies as well as the (remaining) probability that it is assigned to traffic policing. An organization’s team structure is a probability measure $\lambda$ over assignments (so $\lambda\in\Delta\left(\Delta(\mathcal{T})\right)$), together with a reshuffling rate $r_{\nu}$ for each $\nu\in{\mathrm{supp}}(\lambda)$. Thus, a team structure indicates the fraction of the population that gets each type of assignment.
@@
The team structure must respect some basic constraints that are exogenous and indicate the fraction of the organization’s teams that must be assigned to a given task at any given instant, represented by $q_{t}$, where $\sum_{t\in\mathcal{T}}q_{t}=1$. In other words, the organization’s team structure $\lambda$ must satisfy
@@
The analysis from our previous sections was such that there was just one task $t$ with $q_{t}=1$, all teams were assigned to $\nu$ with $\nu(t)=1$, and the designer’s only decision was the reshuffling rate $r_{\nu}$. We are now considering the more general case in which there are several tasks and discretion over how teams are assigned to those tasks. For example, a team structure of a police organization might have 40% of its teams assigned to only do the task of traffic policing and such that those teams are reshuffled after every period; the remaining 60% of teams are assigned half-time to the task of responding to emergencies and half-time to the task of traffic policing, and the teams with mixed assignments are never reshuffled. Together, this would lead to 70% of the population in teams that are traffic policing at any point in time and 30% of the population in teams that are responding to emergencies.
@@
Indeed, when there is more than one task, it can be optimal to have some teams mix across tasks and others constantly assigned to the same task. The essential insights are easy to see in the case of just two tasks, and so we work with that case, i.e., $\mathcal{T}=\{t_{G},t_{B}\}$ with corresponding fractions $q_{G},q_{B}$ of teams that must cover each task (we write $q_{G},q_{B}$ instead of $q_{t_{G}},q_{t_{B}}$ for ease of notation). For simplicity, we further assume that task $t_{G}$ has an arrival probability $a_{G}>0$ of the good game $G$ and no probability of $B$, and correspondingly, task $t_{B}$ has an arrival probability $a_{B}>0$ of the good game $B$ and no probability of $G$. We therefore refer to tasks $t_{B}$ and $t_{G}$ as “bad” and “good” tasks, respectively. This is not to say that completing the bad task is harmful to the organization; in fact, a share $q_{B}$ of teams must be covering that task. The “bad” simply refers to the fact that cooperation, compared to noncooperation on that task, is worse for society. For example, even though traffic policing might present frequent opportunities to accept bribes, traffic policing is valuable to society and needs to be done, and so there is an exogenous organizational constraint indicating the share of the population that needs to be assigned to it.
@@
An organization with only generalists—i.e., where every team faces the same set of tasks—is captured by $\nu^{gen}(t_{G})=q_{G}$ and $\nu^{gen}(t_{B})=q_{B}$ and $\lambda(\nu^{gen})=1$. Another team structure is full specialization, with one assignment being to only the good task, $\nu^{G}(t_{G})=1$, and another assignment being to only the bad task, $\nu^{B}(t_{B})=1$, with $\lambda(\nu^{G})=q_{G}$ and $\lambda(\nu^{B})=q_{B}$. In this case, the designer optimally keeps teams on the good task together ($r_{G}=0$; again, we write $r_{G}$ instead of $r_{\nu^{G}}$ for simplicity, as we do for $r_{B}$) and reshuffles those on the bad task ($r_{B}=1$). This specialization of teams overcomes two of the three conditions required for optimal cooperation from Proposition 1 (conditions 1 and 3): First, greater temptation to deviate in good than bad games still means that desirable cooperation leads to collusion; however, the only teams that encounter bad games are frequently reshuffled and hence never cooperate. Second, good and bad games are fully separated; hence, there are no spillovers constraining the relative frequencies of good and bad games. Also, because cooperation becomes easier when interactions occur more frequently, the condition for optimal cooperation with full specialization on good games is weaker than its analog (condition 2). Thus, full specialization is an optimal team structure if cooperation can be sustained on the good task by itself. This is easy to see, and so we simply state it as an observation.

### v3

@@ Abstract
Teamwork is vital in many settings, and it is socially beneficial for teams to cooperate in some situations (“good games”) and not in others (“bad games;” e.g., those that allow for corruption). A team’s cooperation in any given game depends on expectations of cooperation in future iterations of both good and bad games. We identify when sustaining cooperation on good games necessitates cooperation on bad games. We then characterize how a designer should optimally assign workers to teams and teams to tasks that involve varying arrival rates of good and bad games. Our results show how organizational design can be used to promote cooperation while minimizing corruption.
@@ 1 Introduction
Teamwork—informal cooperation among members—is essential in almost all forms of production, both private and public. However, team members can cooperate in ways that are harmful to a firm, organization, or society. Importantly, “good” and “bad” forms of cooperation are intertwined. For example, building a cooperative culture among police not only induces productive cooperation, such as aiding other officers in dangerous circumstances, but can also lead them to cooperate in corrupt or destructive ways, such as the “blue wall of silence” (police officers under-reporting misconduct). [^1] See Chevigny (1995). Similarly, some scholars (e.g., Gibson and Singh (2003)) document a “white wall of silence”: doctors not reporting colleagues’ malpractice. If a police officer does not collude with a partner who has taken a bribe or used excessive force, that partner may not back the officer up in a future dangerous situation. [^2] Famously, in 1970, Frank Serpico, then a police detective in the NYPD, exposed police corruption by contributing to a New York Times article. A year later, after being shot and wounded during a drug raid, Serpico alleged that his police partners led him into a dangerous situation to be murdered (Phalen, 2012). Ostracism of whistleblowers is anticipated and widespread: “Officers who report misconduct are ostracized and harassed; become targets of complaints and even physical threats; and are made to fear that they will be left alone on the streets in a time of crisis. This draconian enforcement of the code of silence fuels corruption because it makes corrupt cops feel protected and invulnerable” (Mollen Commission (1994)). Similar interactions between good and bad forms of cooperation arise in a wide variety of organizations, including in government, military, hospitals, consulting, construction, manufacturing, and retail.
@@
Our analysis takes a design perspective. By adjusting team composition and assignments to various tasks, an organizational designer can control the incentives and behavior of the agents involved without observing actions in specific games or even whether a game occurs. This opens a new set of mechanism design problems, in which the dynamic set faced is controlled by the designer. This design perspective involves several dimensions.
@@
We also allow the assignment of teams to tasks to be dynamic and reactive—i.e., the current task assignment can be a function of past assignments. For example, a team may be rotating between staying on a task with only good games (for some time) and staying on a task with only bad games (for some time). Or a team may be randomly moved between tasks. We also consider the case where the designer observes whether a game occurred (but not the equilibrium play) and can condition future task assignments on that. The space of such “reactive assignments” is large, and so we focus on a special case where cooperation on bad games is easy to sustain. In an optimal assignment, when the designer observes whether a good game occurred, teams are assigned to the task with more good games until a good game occurs, at which point they are reassigned to tasks with mostly bad games for a number of periods before returning. We derive a similar result for the case where the designer does not observe good games: the difference is that teams move on from the task with good games whether or not a good game occurred.
In the online appendix, we provide comparative static results with regard to the payoff structure. If an organization can directly affect payoffs, which is not our main focus, then these comparative statics shed light on other strategies for enabling beneficial cooperation. A designer can encourage good cooperation while deterring bad cooperation by altering the benefits of cooperation in good games (for instance, giving bonuses for overall profits) or increasing the temptations of defecting in bad games (for instance, offering whistleblower rewards). Interestingly, these behave quite differently than in standard repeated games, where only the ratio of benefits to temptations matters. With multiple games, benefits from cooperation in a given game have spillovers that enable cooperation in all games, while temptations to deviate affect only one game’s incentives at a time. Thus, the latter is more selective than the former. The effective costs of altering the benefits from cooperation or temptations of defecting also vary. Because bonuses are paid, a policy of giving bonuses for cooperative behavior imposes an actual cost on an organization. In contrast, whistleblower rewards deter collusion and are not dispersed on the equilibrium path.
@@ 2 The Model
Properties 13 are standard: Property 1 states that a player’s payoff is only a function of their own action and the aggregate actions, not their own identity and the identities of others, Property 2 implies that players benefit from cooperation, while Property 3 means each player individually benefits from not cooperating in a single game. Together with Property 4, Property 3 also implies a lower bound on the payoffs a player receives in any continuation, as well as an equilibrium that achieves them. Lastly, Properties 5 and 6 simplify the analysis by ensuring that, on the equilibrium path, asymmetric play is never player-optimal, i.e., in any particular game, there is either full cooperation or full non-cooperation. As we show in Lemma 1, these properties allow us to characterize the games using just a few parameters described below. Thus, we can focus our attention on interactions across games, rather than the details of dynamics within a game, which can be quite complex absent such structure. [^4] For example, see Olszewski and Safronov (2018). Beyond the dynamic differences, their game also involves private information.
@@
Other games, such as a favor exchange game in which players randomly have chances to either give or receive favors—both good and bad (e.g., lending money vs. not reporting a friend’s crime)—also satisfy these properties. Different players move at different points in time, but a similar analysis applies.
@@ 2.2 Player-Optimal Equilibria
As there can be multiple equilibria (the usual folk-theorem arguments apply), we focus on those that are optimal for the players. Given the game’s symmetry and its required properties, we characterize the equilibria that maximize the sum of players’ expected discounted payoffs.
Characterizing such player-optimal equilibria makes sense for three reasons. First, they are extremal—supporting maximal cooperation—and thus define the frontier and thus the full set of equilibrium payoffs, regardless of what one believes about equilibrium selection. Second, it may not be robust to presume that a designer can guide players to one equilibrium if all players prefer another. [^5] Our equilibrium selection coincides with that in Rotemberg and Saloner (1986), who focus on the most collusive symmetric equilibrium in a stochastic repeated game. Third, the possibility of interaction across good and bad games that is highlighted here is the object of interest, and (Pareto-dominated) equilibria that involve no such interaction could also be analyzed, but involve no new insights or intuitions. More generally, once one departs from the player-optimal equilibrium, equilibrium behavior need not admit simple characterizations, as we show below, and the designer’s problem no longer reduces to transparent incentive constraints. Analyzing arbitrary equilibria would, therefore, require additional structure or selection assumptions and would obscure the mechanism that is central to our analysis.
@@
Visually depicting the different cases of equilibrium cooperation characterized in Lemma 1 is useful. Figure 1 shows which case occurs as a function of the probabilities of good, resp. bad, games occurring times $\delta/(1-\delta)$. Taking Inequality (1) as an equality delineates the region where total cooperation prevails. Cooperation only in some game $g$ precludes total cooperation, so the former region must be southwest of the latter. For cooperation to be sustainable, Inequality (2) must hold; its equality delineates the region with (some) cooperation. Note: there cannot be cooperation only in game $g$ if the temptation to deviate in that game exceeds the deviation temptation in the other game; thus, only one such region exists.
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{B}/c_{B}$region 2region 1region 3
@@
$\frac{\delta}{1-\delta}p_{B}$$\frac{\delta}{1-\delta}p_{G}$$\max\{{d_{g}}\}/c_{G}$$\max\{{d_{g}}\}/c_{B}$$d_{G}/c_{G}$region 2region 1region 4
@@
Although the games are independent, their strategic interactions necessitate a holistic analysis to understand players’ actions in one of them. For example, myopically improving the incentives for desirable actions in one type of situation may backfire, generating detrimental behaviors elsewhere. Similarly, collusive behavior may be minimized by eradicating situations where it occurs (decreasing $p_{B}$), reducing its benefits (decreasing $c_{B}$), or incentivizing team members to stop colluding or to report it (increasing $d_{B}$); although all these may serve their intended purpose of reducing collusion, they also can impair productive cooperation in other situations if incentives elsewhere are not adjusted to compensate. The dual of this warning highlights missed opportunities: A myopic design approach may overlook opportunities to reduce bad cooperation by overlooking opportunities to reduce the need for good cooperation. We discuss the implications of changing payoffs in greater detail in the online appendix.
@@ 3 How Cooperation Depends on Team Durability
We model the reshuffling of teams (reducing team durability) as a probability $r$ with which the stochastic game of a team ends after each period. Given the continuum of workers, we assume that reshuffled players never face a past teammate, that there is no spreading of information, and that players only condition on histories with their current team. [^7] This greatly simplifies the analysis; see, e.g., Ali and Miller (2016b) for a treatment where such contagion is possible. Our simplification also requires sufficiently many workers so that completely new teams can be formed, as well as anonymity of workers outside their teams (deviations can be punished only by current team members). These conditions are more likely to hold—at least approximately—in large organizations, potentially leading to differences by organization size that we leave for future research. This can also be handled by hiring people part-time or by outsourcing and subcontracting tasks on short-term rotations. Consequently, a reshuffling rate $r$ is equivalent to changing the discount factor $\delta$ to $(1-r)\delta$; essentially, the designer can induce any effective discount factor of at most $\delta$ by varying team durability. Thus, formally, we characterize optimal reshuffling by examining the player optimal equilibria for a given team (with histories only applying to that team) as a function of discount factor $(1-r)\delta$.
@@
The temptation to deviate in good games must be no higher than that in bad games; otherwise, whenever good cooperation is possible, then so is bad. Thus, condition 1 is necessary. Cooperation being sustainable in the good game by itself (condition 2) is also necessary because, for optimal cooperation, there are no strategic interdependencies across the games. Lastly, good games must have a sufficiently high probability, relative to bad games (condition 3), to make it feasible (possibly by reshuffling) to induce players to cooperate in the good game but not in the bad one: it ensures that there exists an effective discount rate $(1-r)\delta$ such that Equation (1) fails, while (2) holds for the good game. The proposition then follows from a careful application of Lemma 1.
@@ 4 Assigning Teams to Tasks
In the previous sections, the arrival rates of good and bad games were fixed, and the only design dimension was the rate at which teams were reshuffled. We now consider situations in which there are different tasks to which teams can be assigned, and the arrival rates of the two games differ across these tasks, thereby allowing an organizational designer to control arrival rates. For example, in a police force, the task of responding to emergencies might create infrequent but important situations in which backup is required (good games), whereas the task of traffic policing might present frequent opportunities to accept small bribes (bad games).
As a running illustration of optimal assignments in this and the next section, consider the following example.
Example 1.
Payoffs are such that $\delta=3/5,c_{G}=c_{B}=1,d_{G}=1$, and $d_{B}=1/2$, and the arrival probabilities of games are $p_{G}=1/4$ and $p_{B}=3/8$.
In this setting, cooperation on good games is not sustainable. To see this, note that (1) is violated. That is, the payoff from deviating in a good game and not cooperating exceeds the maximum potential future payoff loss, presuming that there could be full cooperation on both games:
$d_{G}=1>\frac{15}{16}=\frac{\frac{3}{5}}{1-\frac{3}{5}}\cdot\left(\frac{1}{4}\cdot 1+\frac{3}{8}\cdot 1\right)=\frac{\delta}{1-\delta}(p_{G}c_{G}+p_{B}c_{B}).$
Suppose, however, that the relative arrival rates of good and bad games can be adjusted. In particular, there are two tasks that have to be covered with equal weights—at any point in time, half of the teams need to cover each—and on one task, (only) good games arrive at a rate of $a_{G}=1/2$, and on the other task (only) bad games arrive at a rate of $a_{B}=3/4$. If a team is randomly assigned to each task in each period with equal probability, then this satisfies the task coverage constraint and gives back $p_{G}=1/4=\frac{1}{2}a_{G}$ and $p_{B}=3/8=\frac{1}{2}a_{B}$ as the arrival probabilities of future games.
But now, instead of assigning a team to each task with equal probability, one could assign it to different mixtures of tasks. Ideally, assigning it entirely to the task with good games would sustain cooperation. Then, other teams could be assigned to the task with bad games (so that each task is covered) and reshuffled to preclude cooperation, leading to optimal cooperation.
However, note that in this example, if one assigns a team entirely to the task with only good games, that would also not lead to cooperation on good games since future good games alone do not provide enough incentives to cooperate as the incentives to deviate exceed any benefits from future cooperation:
$d_{G}=1>\frac{3}{4}=\frac{\frac{3}{5}}{1-\frac{3}{5}}\cdot\frac{1}{2}\cdot 1=\frac{\delta}{1-\delta}a_{G}c_{G}$
(a condition which we later formalize in Observation 1).
Thus, cooperation on good games cannot be sustained with the original default arrival rates, nor with assignment to just the task with only good games. Instead, what is necessary to sustain cooperation is an assignment that includes a sufficient stream of future bad games that then provide rewards and incentives not to deviate in a current good game. Indeed, there is cooperation on both games when teams face a high enough share of future bad games since $\max_{g}\{d_{g}\}=1<\frac{9}{8}=\frac{\frac{3}{5}}{1-\frac{3}{5}}\cdot\frac{3}{4}\cdot 1=\frac{\delta}{1-\delta}a_{B}c_{B}$.
To illustrate, consider a task assignment in which a team is assigned to the task with only good games with probability $\nu_{G}$ and to the task with only bad games with probability $\nu_{B}=1-\nu_{G}$. The assignment that provides enough incentives to sustain cooperation in good games, while minimizing the share of bad games, solves
$\max\left\{\frac{1}{2},1\right\}=\frac{\frac{3}{5}}{1-\frac{3}{5}}\left(\nu_{G}\cdot\frac{1}{2}\cdot 1+\nu_{B}\cdot\frac{3}{4}\cdot 1\right),$ (3)
and is thus given by $\nu_{G}=1/3$ and $\nu_{B}=2/3$.
What Example 1 shows is that if the organizational designer has additional flexibility to steer a team towards tasks that involve more good or bad games, then they can find mixes that sustain cooperation with minimal arrivals of bad games.
The example illustrates just one part of the overall design problem. The designer still has to cover all tasks with teams, that is, in the previous example, achieve the 50–50 assignment of teams to tasks on average. So, by tilting some teams toward this optimal mix, some tasks will not be fully covered. Those can be covered with other teams, but ones in which the designer simply gives up on motivating cooperation. As a result, some teams end up incentivized to cooperate, while others are left without cooperation—the assignment will often be asymmetric. Exactly which assignment ends up being optimal depends on the setting: the relative arrival rates, payoffs, and the required task coverage.
@@
We model this additional design option as follows. There is a finite set of tasks $\mathcal{T}$. Each task $t$ has its own probability of good and bad games, $p_{G}(t)$ and $p_{B}(t)$. An assignment for a team, denoted $\nu$, is a probability measure on $\mathcal{T}$ that specifies the probability with which the team is assigned to each task, with the realized task redrawn in every period. For example, if there are two tasks, responding to emergencies vs. traffic policing, then $\nu$ indicates the probability that the team is assigned to respond to emergencies as well as the (remaining) probability that it is assigned to traffic policing.
An organization’s team structure is a probability measure $\lambda$ over assignments (so $\lambda\in\Delta\left(\Delta(\mathcal{T})\right)$), together with a reshuffling rate $r_{\nu}$ for each $\nu\in{\mathrm{supp}}(\lambda)$. Thus, a team structure indicates the fraction of the population that gets each type of assignment.
@@
The team structure must respect an exogenous constraint that specifies the fraction of the organization’s teams assigned to a given task at any given instant, represented by $q_{t}$, where $\sum_{t\in\mathcal{T}}q_{t}=1$. In other words, the organization’s team structure $\lambda$ must satisfy
@@
For reactive assignment $\nu$, the average arrival probabilities of the different games are given by $p_{g}(\nu)\coloneqq\sum_{t\in\mathcal{T}}\nu(t)\,p_{g}(t)$. For $p_{G}(\nu),p_{B}(\nu)>0$, equilibrium play is still characterized Lemma 1, that is, there are still the four types of canonical cooperation, and, of course, when only one game has a positive probability of arrival, then cooperation is sustained if and only if $d_{G}\leq\frac{\delta}{1-\delta}p_{g}(\nu)c_{g}$. This implies a well-defined social value. For any team structure $\lambda$, together with reshuffling rate $r_{\nu}$ for each $\nu\in{\mathrm{supp}}(\lambda)$, let $\chi_{g}(\nu)\in\{0,1\}$ denote the equilibrium indicator of cooperation in game $g\in\{G,B\}$ under assignment $\nu$. The social value generated by team structure $\lambda$ is then
$\sum_{\nu\in{\mathrm{supp}}(\lambda)}\lambda(\nu)\sum_{g\in\{G,B\}}p_{g}(\nu)\,\chi_{g}(\nu)\,V_{g}.$
This can be taken as an objective function for the designer in choosing $\lambda$.
@@
This model nests the model from our previous sections, where there was just one task $t$ with $q_{t}=1$, all teams were assigned to $\nu$ with $\nu(t)=1$, and the designer’s only decision was the reshuffling rate $r_{\nu}$. We are now considering the more general case in which there are several tasks and discretion over how teams are assigned to those tasks. [^8] The previous model is also nested if interpreted as there being several tasks, but the designer has no discretion over task assignments; i.e., they must use the same assignment for each team, namely $\nu$ with $\nu(t)=q_{t}$.
@@
The essential insights are easy to see in the case of just two tasks, as in Example 1, and so we work with that case. Specifically, let $\mathcal{T}=\{t_{G},t_{B}\}$ with corresponding fractions $q_{G},q_{B}$ of teams that must cover each task (we write $q_{G},q_{B}$ instead of $q_{t_{G}},q_{t_{B}}$ for ease of notation). For simplicity, we further assume that task $t_{G}$ has an arrival probability $a_{G}=p_{G}(t_{G})>0$ of the good game $G$ and no probability of $B$, and correspondingly, task $t_{B}$ has an arrival probability $a_{B}=p_{B}(t_{B})>0$ of the bad game $B$ and no probability of GG. [^9] Nothing substantive changes with tasks that are not extremal; i.e., when they generate both good and bad games. The designer has less flexibility in shaping the stream of games faced by a team, since assignments cannot fully separate the two types. Similarly, introducing more than two tasks with various heterogeneous arrival rates complicates the full characterization of the optimal assignment, but not the basic forces that we analyze of trading off more social and less socially desirable arrival rates to generate cooperation with minimal relative amounts of cooperation in bad games. We therefore refer to tasks $t_{G}$ and $t_{B}$ as “good” and “bad” tasks, respectively. This is not to say that completing the bad task is harmful to the organization; in fact, a share $q_{B}$ of teams must be covering that task. The “bad” simply refers to the fact that cooperation, compared to noncooperation on that task, is worse for society. For example, even though traffic policing might present frequent opportunities to accept bribes, it is valuable to society and must be done, so there is an exogenous organizational constraint specifying the share of the population that must be assigned to it.
@@
An organization with only generalists—i.e., where every team faces the same set of tasks—is captured by a single reactive assignment $\nu^{gen}$ with $\nu^{gen}(t_{G})=q_{G}$ and $\nu^{gen}(t_{B})=q_{B}$. Another team structure is full specialization, with one assignment being to only the good task, $\nu^{G}(t_{G})=1$, and another assignment being to only the bad task, $\nu^{B}(t_{B})=1$ (task coverage then requires $\lambda(\nu^{G})=q_{G}$ and $\lambda(\nu^{B})=q_{B}$). In this case, the designer optimally keeps teams on the good task together ($r_{G}=0$; again, we write $r_{G}$ instead of $r_{\nu^{G}}$ for simplicity, as we do for $r_{B}$) and reshuffles those on the bad task ($r_{B}=1$). This specialization of teams overcomes two of the three conditions required for optimal cooperation from Proposition 1 (conditions 1 and 3): First, greater temptation to deviate in good than bad games still means that desirable cooperation leads to collusion; however, the only teams that encounter bad games are frequently reshuffled and hence never cooperate. Second, good and bad games are fully separated; hence, there are no spillovers constraining the relative frequencies of good and bad games. Also, because cooperation becomes easier when interactions occur more frequently, the condition for optimal cooperation with full specialization on good games is weaker than its analog (condition 2). Thus, full specialization is an optimal team structure if cooperation can be sustained on the good task by itself. This is easy to see, and so we simply state it as an observation.